The percent of N in 66% pure (NH402 4 sample isKkkkkkk

Grade:10

3 Answers

Vikas TU
6 years ago
I think your compound is (nh4)2 so4 
presently, 
atomic mass of compound=2*(14+1*4)+32+16*4 
=2*18+32+64 
= 36+96 =132g 
nuclear weight of nitrogen=28g 
along these lines, 
% of nitrogen in 66% =28/(66/100 *132) %=32 % approx 
Trisha Barui
6 years ago
atomic wt= 2*(14+1*4)+32+16*4 =132
so % of nitrogen = 28/132*100=21.2%
so % of nitrogen in 66% is 67/100*21.2=14%
 
homo sapiens
6 years ago
The molar mass of (NH​4​)​2SO4​   is  =132g
now,
  if we have,
        ​28g of N​2​ -------------------------- in  ​132g of (NH​4)2SO4
​       then we have,
​        x% of N2​---------------------------- in 100%  of |---|------|
         therefore,
​                  28 * 100/132   =21.2121212121...  =21.2​%
​                 now,
​                       if we have,
​                   21.2% of N2----------------------------------------in 100% 0f that solution
​                   then,we have,
​                   x% of N2​--------------------------------------------in 66% of    that solution
            
​               therefore,
              
​                 21.2 * 66 / 100 = 13.992  =14
    therefore  we have 14% of N2​ in 66% pure (NH4)​2SO4.